OpenStudy (anonymous):

When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the first equation? 9 - x y - 9/2 9 - 2y 1/2 y + 9

OpenStudy (anonymous):

9-2y

Directrix (directrix):

Second Equation: x - y = 6 Substitute for x in the first so solve the second equation for x. x - y = 6 x = 6 + y I don't see 6 + y as an option.

OpenStudy (anonymous):

im 100 percent sure its c

Directrix (directrix):

I am thinking that 9 - 2y is the answer to the following question: When solving the system x - y = 6 and x + 2y = 9, which expression could be substituted for x in the first equation? I am wondering if the question is posted correctly.